OpenStudy (anonymous):

Write the equation for the ellipse with horizontal minor axis of length 8 and foci located at (0,5) and (0,-5).

OpenStudy (dumbcow):

length of major axis = 2a minor axis = 2b length of foci = c such that ..... b^2 +c^2 = a^2 from the info given : b = 4 c = 5 therefore, 16 +25 = a^2 --> a^2 = 41

OpenStudy (anonymous):

so then x^2/16+y^2/41=1

OpenStudy (anonymous):

oh wait nope.. a is the x-int's so wouldn't it be x^2/41+y^2/16=1

OpenStudy (dumbcow):

sorry i was away for a moment ..... "a" refers to major axis major axis is vertical or "y-axis" so your 1st answer is correct x^2/16 + y^2/41 = 1